OpenStudy (anonymous):

sqrt (5x+21) = x+3solve

OpenStudy (anonymous):

5x+21 = x^2 + 6x +9 x^2 + x - 12 = 0 (x + 4)(x - 3) = 0 therefore x = 3 or -4

jhonyy9 (jhonyy9):

sqrt(5x+21)=x+3 squared both sides 5x+21=(x+3)(x+3) 5x+21=x2+6x+9 x2 +x -12=0 (x+4)(x-3)=0 x+4=0 x=-4 and x-3=0 x=3 x_1=-4 x_2=3